A car is traveling at 9 m/s in a circle that has a radius of 60 m. what must be the minimum value of ΞΌ for the car to make the turn without skidding?


1
Expert's answer
2021-11-16T13:40:57-0500
"F_c=F_{fr},""\dfrac{mv^2}{R}=\mu mg,""\mu=\dfrac{v^2}{Rg},""\mu=\dfrac{(9\ \dfrac{m}{s})^2}{60\ m\times9.8\ \dfrac{m}{s^2}}=0.14"